If 4 boys can shovel a driveway in 2 hours, how many minutes will 5 boys take to do the job?
Answer by nyc_function(2733)


Since 4 boys can shovel a driveway in 2 hours, 1 boy can do it in 4 x 2 hours = 8 hours. Thus, 5 boys can do it in 8/5 hours. Now, each hour is 60 minutes, so 5 boys can do it in 8/5 * 60 = 60 + 36 = 96 minutes.
How about now?